OpenStudy (anonymous):

please help me i've tried this a bunch of times and keeping the wrong answer. equation below

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?} xy dA\] R is bounded by x=0, y= 2x +1, and y= -2x +5

OpenStudy (anonymous):



OpenStudy (anonymous):

if you solve for where they intersect then it is 1

OpenStudy (anonymous):

x=1 y= 2x +1 y=2*1+1 y=3

OpenStudy (anonymous):

that is your bounds

OpenStudy (anonymous):